OpenStudy (anonymous):

use double integral to find volume of tetrahedron bounded by the cooridnate planes and the plane 5x+5y+2z=10

OpenStudy (anonymous):

i'm down to do this whole thing step by step. first thing: do you know how to graph the plane?

OpenStudy (anonymous):

Lol yeah I was about to say that I think what would help the best is to first graph the picture

OpenStudy (anonymous):

I can do the int without much issue i need a little direction setting up the problem.

OpenStudy (anonymous):

Euler seems to be more confident about teaching you it so I will let him expain it to you since I'm actually studying to take a final for this in calc 3 and although this would be good review, I think you definitely want someone who can easily explain it to you

OpenStudy (anonymous):

you need to graph the boundaries to know the limits of the integrals, which in this case is a plane. do you know how to graph this plane?

OpenStudy (anonymous):

thats where i am a bit fuzzy.

OpenStudy (anonymous):

to graph a plane, all you need to do is find the intercepts. similarly to 2 dimensions, where finding x-intercepts are when y = 0, in 3-d, the x-intercepts are when both y and z = 0. when y = z = 0 --> 5x = 10 . the x-intercept of the plane is x = 2 when x = y = 0 --> 2z = 10. the z-intercept of the plane is z = 5 y intercept = 2 here is the graph:  it's the tetrahedron between origin and the plane. the x-y plane looks like this: 

OpenStudy (anonymous):

so the integrating limits of x are from x = 0 to x = 2 and the integrating limits of y are from y = 0 to y = 2 - x

OpenStudy (anonymous):

now i feel like a idiot....one should never start homework so late at night.

OpenStudy (anonymous):

the function you are integrating is f(x,y) = z = 1/2(5x + 5y) \[1/2 \int\limits_{0}^{2}\int\limits_{0}^{2-x} (10 - 5x - 5y) dy dx\]

OpenStudy (anonymous):

CORRECTION: the function you are integrating is f(x,y) = z = 1/2(10 - 5x - 5y)